OpenStudy (anonymous):

Is my answer correct? Question in integration the question: 1/2X^3 Solution: -1/X^2

OpenStudy (therealmeeeee):

Yes I it is I think let me check

OpenStudy (anonymous):

ok

OpenStudy (therealmeeeee):

yes it right Goodnight!!

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

\[\int\frac{1}{2x^3}dx = \int2x^{-3}dx = 2\frac{x^{-3+1}}{-3 + 1} = -x^{-2} =-\frac{1}{x^2} \]

OpenStudy (anonymous):

yes thanks @bhl6180 my answer correct